A 40t rail car travels at 4km/h and collides with a 100t wagon on the same track, moving in the opposite direction at 1,2km/h

i) Find their velocities immediately after impact, assuming no energy loss [8]

ii) Find the impulse between them [5]


1
Expert's answer
2021-03-01T12:37:55-0500

Given,

Mass of the rail car "(m)=40 ton"

Speed of rail car "(u)=4km\/h"

Mass of wagon "(M)=100 ton"

Speed o the wagon "(U)=1.2 km\/h"

As per the question, both are moving in opposite direction.

Let after the collision, the velocities becomes "v" and "V" respectively.

Now, applying the conservation of energy

"mu-MU=mv+MV"

"\Rightarrow 40\times 4-100\times 1.2=40v+100V"

"\Rightarrow 40v+100V=160-120"

"\Rightarrow 40v+100V=40"

"\Rightarrow 2v+5V=2...(i)"

As there is no loss of energy, so total energy always be conserve,

So, the coefficient of elasticity "(e)=1"

"e=\frac{V-v}{u-U}"

"\Rightarrow V-v=4-1.2"

"\Rightarrow V-v=2.8km/h ...(ii)"

From equation (i) and (ii)

"\Rightarrow 7V=5.6+2"

"\Rightarrow V=\frac{7.6}{7}km/hour"

"V=1.08km\/hour"

Now, substituting the value of V in equation (ii)

"v=-2.8+1.08"

"=-1.72 km\/hour"

here -ive sign indicates that after the collision, both car will move opposite to each other.

ii) Impulse between them = "(mu+mv)+(MU+MV)"

"=40\times 4+40\times 1.72+100\times 1.2+100\times 1.08"

"=160+68.8+120+108"

"=456.8ton-km\/hour"